OpenStudy (anonymous):

divide and answer in a+b¡ form. and how to Check the answer by using multiplication?? 2/1-√-4

OpenStudy (anonymous):

2/(1-2i) = 2(1+2i)/(5)

OpenStudy (dumbcow):

\[\sqrt{-4} = i \sqrt{4} = 2i\] \[\frac{2}{1-2i}*\frac{1+2i}{1+2i} = \frac{2(1+2i)}{5}\]

OpenStudy (dumbcow):

i^2 = -1